OpenStudy (anonymous):

Is this trinomial a trinomial square? x^2-12x+36

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Don't just give out answers, explain yourself.

OpenStudy (anonymous):

Yeah, how is this a trinomial square?

OpenStudy (anonymous):

x^2-12x+36=(x-6_*(x-6)=(x-6)^2

OpenStudy (anonymous):

the underscore is supposed to be a ) Typing error.

OpenStudy (anonymous):

Thank you for explaining.

OpenStudy (anonymous):

Is this a trinomial square? 9x^2+81x+9

OpenStudy (anonymous):

@kathyzhu23

OpenStudy (anonymous):

I'm not too sure about that one, I'm sorry